# EMMA-500: Enhancing Massively Multilingual Adaptation of Large Language Models

### v1

@@ D.1 Commonsense Reasoning
Llama 2 7B 0.7247 0.8791 0.0068 0.6627 0.0522 0.7028 0.0148 0.7224 0.0277 0.6825 0.0263 0.6984 0.0205
Llama 2 7B Chat 0.6945 0.8555 0.0073 0.7108 0.0501 0.6840 0.0150 0.6464 0.0295 0.6571 0.0268 0.6131 0.0217
CodeLlama 2 7B 0.7092 0.8452 0.0075 0.6747 0.0517 0.6361 0.0155 0.7224 0.0277 0.6667 0.0266 0.7103 0.0202
LLaMAX Llama 2 7B 0.6749 0.7789 0.0086 0.6145 0.0537 0.7101 0.0147 0.6540 0.0294 0.5556 0.0280 0.7361 0.0197
LLaMAX Llama 2 7B Alpaca 0.6986 0.8275 0.0078 0.6627 0.0522 0.7237 0.0144 0.6616 0.0292 0.5937 0.0277 0.7222 0.0200
MaLA-500 Llama 2 10B v1 0.6589 0.8366 0.0077 0.6386 0.0531 0.6017 0.0158 0.6920 0.0285 0.6190 0.0274 0.5655 0.0221
MaLA-500 Llama 2 10B v2 0.6589 0.8366 0.0077 0.6386 0.0531 0.6017 0.0158 0.6920 0.0285 0.6190 0.0274 0.5655 0.0221
YaYi Llama 2 7B 0.7450 0.8852 0.0066 0.7108 0.0501 0.7226 0.0145 0.7414 0.0270 0.7175 0.0254 0.6925 0.0206
TowerBase Llama 2 7B 0.7429 0.8714 0.0069 0.7470 0.0480 0.6945 0.0149 0.7567 0.0265 0.6476 0.0270 0.7401 0.0196
TowerInstruct Llama 2 7B 0.7400 0.8628 0.0071 0.7711 0.0464 0.6840 0.0150 0.7719 0.0259 0.6635 0.0267 0.6865 0.0207
Occiglot Mistral 7B v0.1 0.7461 0.8654 0.0071 0.7952 0.0446 0.6601 0.0153 0.7186 0.0278 0.6635 0.0267 0.7738 0.0187
Occiglot Mistral 7B v0.1 Instruct 0.7293 0.8589 0.0072 0.7470 0.0480 0.6455 0.0155 0.7072 0.0281 0.6476 0.0270 0.7698 0.0188
BLOOM 7B 0.7013 0.8224 0.0079 0.7108 0.0501 0.5881 0.0159 0.7681 0.0261 0.5746 0.0279 0.7440 0.0195
BLOOMZ 7B 0.6795 0.8340 0.0077 0.7349 0.0487 0.5756 0.0160 0.6730 0.0290 0.5492 0.0281 0.7103 0.0202
mGPT 0.5969 0.6267 0.0100 0.5904 0.0543 0.5349 0.0161 0.5741 0.0305 0.5810 0.0278 0.6746 0.0209
mGPT 13B 0.6359 0.7062 0.0094 0.6386 0.0531 0.5777 0.0160 0.6122 0.0301 0.6000 0.0276 0.6806 0.0208
XGLM 7.5B 0.6884 0.7940 0.0084 0.6506 0.0527 0.6496 0.0154 0.6768 0.0289 0.6349 0.0272 0.7242 0.0199
YaYi 7B 0.6979 0.8404 0.0076 0.7590 0.0472 0.5808 0.0159 0.7224 0.0277 0.5587 0.0280 0.7262 0.0199
Llama 3 8B 0.7684 0.8680 0.0070 0.7108 0.0501 0.7529 0.0139 0.7985 0.0248 0.7143 0.0255 0.7659 0.0189
Llama 3.1 8B 0.7552 0.8757 0.0068 0.6627 0.0522 0.7529 0.0139 0.8061 0.0244 0.6762 0.0264 0.7579 0.0191
Gemma 7B 0.7741 0.8800 0.0067 0.7349 0.0487 0.7539 0.0139 0.7833 0.0255 0.7206 0.0253 0.7718 0.0187
Gemma 2 9B 0.8007 0.8942 0.0064 0.7590 0.0472 0.7977 0.0130 0.8289 0.0233 0.7429 0.0247 0.7817 0.0184
Qwen 2 7B 0.7644 0.8632 0.0071 0.6988 0.0507 0.6882 0.0150 0.7871 0.0253 0.7175 0.0254 0.8313 0.0167
Qwen 1.5 7B 0.7259 0.8331 0.0077 0.6988 0.0507 0.6590 0.0153 0.7034 0.0282 0.6635 0.0267 0.7976 0.0179
EMMA-500 Llama 2 7B 0.7280 0.8245 0.0079 0.7229 0.0494 0.7164 0.0146 0.6920 0.0285 0.6921 0.0261 0.7202 0.0200
D.2 Code Generation
@@
D.3 Natural Language Inference
Table 27: 0-shot results (ACC) on XNLI in all languages.
@@ D.3 Natural Language Inference
Llama 2 7B 0.4019 0.3542 0.0096 0.4265 0.0099 0.4711 0.0100 0.3667 0.0097 0.5530 0.0100 0.4052 0.0098 0.5008 0.0100 0.3771 0.0097 0.4237 0.0099 0.3494 0.0096 0.3635 0.0096 0.3727 0.0097 0.3361 0.0095 0.3663 0.0097 0.3618 0.0096
Llama 2 7B Chat 0.3858 0.3442 0.0095 0.3707 0.0097 0.4309 0.0099 0.3815 0.0097 0.5024 0.0100 0.3944 0.0098 0.4482 0.0100 0.3578 0.0096 0.4209 0.0099 0.3422 0.0095 0.3349 0.0095 0.3695 0.0097 0.3390 0.0095 0.3811 0.0097 0.3695 0.0097
CodeLlama 2 7B 0.4019 0.3341 0.0095 0.3775 0.0097 0.4723 0.0100 0.3763 0.0097 0.5478 0.0100 0.4438 0.0100 0.4920 0.0100 0.3594 0.0096 0.4606 0.0100 0.3329 0.0094 0.3502 0.0096 0.3859 0.0098 0.3325 0.0094 0.4040 0.0098 0.3594 0.0096
LLaMAX Llama 2 7B 0.4427 0.3378 0.0095 0.4683 0.0100 0.4896 0.0100 0.4257 0.0099 0.5490 0.0100 0.4759 0.0100 0.4779 0.0100 0.4550 0.0100 0.4554 0.0100 0.4305 0.0099 0.4185 0.0099 0.4329 0.0099 0.4418 0.0100 0.4386 0.0099 0.3438 0.0095
LLaMAX Llama 2 7B Alpaca 0.4509 0.3442 0.0095 0.4639 0.0100 0.4976 0.0100 0.4341 0.0099 0.5811 0.0099 0.4896 0.0100 0.5197 0.0100 0.4562 0.0100 0.4627 0.0100 0.4357 0.0099 0.4080 0.0099 0.4386 0.0099 0.4430 0.0100 0.4309 0.0099 0.3578 0.0096
MaLA-500 Llama 2 10B v1 0.3811 0.3594 0.0096 0.4120 0.0099 0.4751 0.0100 0.3446 0.0095 0.5618 0.0099 0.3410 0.0095 0.4759 0.0100 0.3365 0.0095 0.3394 0.0095 0.3522 0.0096 0.3369 0.0095 0.3382 0.0095 0.3502 0.0096 0.3602 0.0096 0.3325 0.0094
MaLA-500 Llama 2 10B v2 0.3811 0.3594 0.0096 0.4120 0.0099 0.4751 0.0100 0.3446 0.0095 0.5618 0.0099 0.3410 0.0095 0.4759 0.0100 0.3365 0.0095 0.3394 0.0095 0.3522 0.0096 0.3369 0.0095 0.3382 0.0095 0.3502 0.0096 0.3602 0.0096 0.3325 0.0094
YaYi Llama 2 7B 0.4128 0.3414 0.0095 0.4261 0.0099 0.4884 0.0100 0.3735 0.0097 0.5647 0.0099 0.4578 0.0100 0.5104 0.0100 0.3948 0.0098 0.4627 0.0100 0.3570 0.0096 0.3562 0.0096 0.3936 0.0098 0.3349 0.0095 0.3751 0.0097 0.3554 0.0096
TowerBase Llama 2 7B 0.3984 0.3390 0.0095 0.4137 0.0099 0.4787 0.0100 0.3526 0.0096 0.5635 0.0099 0.4169 0.0099 0.4944 0.0100 0.3454 0.0095 0.4594 0.0100 0.3502 0.0096 0.3478 0.0095 0.3574 0.0096 0.3337 0.0095 0.3719 0.0097 0.3522 0.0096
TowerInstruct Llama 2 7B 0.4036 0.3365 0.0095 0.4293 0.0099 0.4884 0.0100 0.3498 0.0096 0.5695 0.0099 0.4651 0.0100 0.4643 0.0100 0.3474 0.0095 0.4627 0.0100 0.3394 0.0095 0.3390 0.0095 0.3787 0.0097 0.3353 0.0095 0.3735 0.0097 0.3747 0.0097
Occiglot Mistral 7B v0.1 0.4235 0.3386 0.0095 0.4137 0.0099 0.5177 0.0100 0.3771 0.0097 0.5586 0.0100 0.5165 0.0100 0.5193 0.0100 0.3574 0.0096 0.4763 0.0100 0.3470 0.0095 0.3739 0.0097 0.4301 0.0099 0.3349 0.0095 0.3863 0.0098 0.4056 0.0098
Occiglot Mistral 7B v0.1 Instruct 0.4081 0.3438 0.0095 0.3884 0.0098 0.5084 0.0100 0.4000 0.0098 0.5566 0.0100 0.4863 0.0100 0.5169 0.0100 0.3430 0.0095 0.4004 0.0098 0.3317 0.0094 0.3635 0.0096 0.3799 0.0097 0.3406 0.0095 0.3759 0.0097 0.3863 0.0098
BLOOM 7B 0.4160 0.3385 0.0067 0.3992 0.0069 0.3978 0.0069 0.3537 0.0068 0.5397 0.0070 0.4882 0.0071 0.4980 0.0071 0.4651 0.0070 0.4303 0.0070 0.3788 0.0069 0.3505 0.0067 0.3509 0.0067 0.4220 0.0070 0.4739 0.0071 0.3535 0.0068
BLOOMZ 7B 0.3713 0.3269 0.0094 0.3402 0.0095 0.4169 0.0099 0.3582 0.0096 0.4687 0.0100 0.3602 0.0096 0.4305 0.0099 0.4040 0.0098 0.3747 0.0097 0.3361 0.0095 0.3309 0.0094 0.3373 0.0095 0.3683 0.0097 0.3667 0.0097 0.3502 0.0096
mGPT 0.4051 0.3382 0.0095 0.4173 0.0099 0.4478 0.0100 0.3542 0.0096 0.4936 0.0100 0.4281 0.0099 0.4406 0.0100 0.4133 0.0099 0.4321 0.0099 0.4189 0.0099 0.3578 0.0096 0.4012 0.0098 0.3422 0.0095 0.4550 0.0100 0.3361 0.0095
XGLM 7.5B 0.4375 0.3349 0.0095 0.4365 0.0099 0.4755 0.0100 0.4040 0.0098 0.5309 0.0100 0.4707 0.0100 0.4542 0.0100 0.4574 0.0100 0.4598 0.0100 0.4570 0.0100 0.4137 0.0099 0.4679 0.0100 0.4193 0.0099 0.4289 0.0099 0.3522 0.0096
YaYi 7B 0.3987 0.3980 0.0098 0.3578 0.0096 0.4261 0.0099 0.3647 0.0096 0.5052 0.0100 0.4771 0.0100 0.4819 0.0100 0.4004 0.0098 0.3912 0.0098 0.3406 0.0095 0.3434 0.0095 0.3317 0.0094 0.3707 0.0097 0.4418 0.0100 0.3494 0.0096
Llama 3 8B 0.4497 0.3365 0.0095 0.4534 0.0100 0.5048 0.0100 0.3928 0.0098 0.5502 0.0100 0.4952 0.0100 0.5056 0.0100 0.4755 0.0100 0.4916 0.0100 0.3892 0.0098 0.4627 0.0100 0.4823 0.0100 0.3349 0.0095 0.4900 0.0100 0.3815 0.0097
Llama 3.1 8B 0.4562 0.3386 0.0095 0.4558 0.0100 0.5145 0.0100 0.3896 0.0098 0.5522 0.0100 0.5028 0.0100 0.5177 0.0100 0.4940 0.0100 0.4916 0.0100 0.3936 0.0098 0.4807 0.0100 0.4932 0.0100 0.3506 0.0096 0.4711 0.0100 0.3976 0.0098
Gemma 7B 0.4258 0.3349 0.0095 0.4349 0.0099 0.4863 0.0100 0.3811 0.0097 0.5205 0.0100 0.4414 0.0100 0.4976 0.0100 0.4434 0.0100 0.4739 0.0100 0.4064 0.0098 0.3795 0.0097 0.4325 0.0099 0.3542 0.0096 0.4329 0.0099 0.3667 0.0097
Gemma 2 9B 0.4674 0.3418 0.0095 0.4952 0.0100 0.5137 0.0100 0.4325 0.0099 0.5345 0.0100 0.5141 0.0100 0.5229 0.0100 0.4731 0.0100 0.4956 0.0100 0.4558 0.0100 0.4988 0.0100 0.5072 0.0100 0.4402 0.0100 0.4570 0.0100 0.3293 0.0094
Qwen 2 7B 0.4277 0.3369 0.0095 0.4546 0.0100 0.4819 0.0100 0.3683 0.0097 0.5426 0.0100 0.4723 0.0100 0.5141 0.0100 0.4498 0.0100 0.4739 0.0100 0.3731 0.0097 0.3827 0.0097 0.4353 0.0099 0.3402 0.0095 0.4357 0.0099 0.3538 0.0096
Qwen 1.5 7B 0.3947 0.3434 0.0095 0.4092 0.0099 0.4277 0.0099 0.3618 0.0096 0.4908 0.0100 0.3787 0.0097 0.4313 0.0099 0.3815 0.0097 0.3888 0.0098 0.3542 0.0096 0.4422 0.0100 0.3884 0.0098 0.3386 0.0095 0.4438 0.0100 0.3398 0.0095
EMMA-500 Llama 2 7B 0.4514 0.3478 0.0095 0.4627 0.0100 0.4707 0.0100 0.4586 0.0100 0.5378 0.0100 0.4707 0.0100 0.4687 0.0100 0.4759 0.0100 0.4655 0.0100 0.4618 0.0100 0.4197 0.0099 0.4486 0.0100 0.4598 0.0100 0.4703 0.0100 0.3522 0.0096
D.4 Math
@@
Tables 29 and 29 show 3-shot results on MGSM by direct and Chain-of-Thought prompting respectively. All the scores are obtained by flexible matching.
@@
Table 28: 3-shot results (ACC) on MGSM by direct prompting and flexible matching.

### v2

@@ Appendix E Detailed Results
Llama 2 7B 72.47 87.91 0.68 66.27 5.22 70.28 1.48 72.24 2.77 68.25 2.63 69.84 2.05
Llama 2 7B Chat 69.45 85.55 0.73 71.08 5.01 68.40 1.50 64.64 2.95 65.71 2.68 61.31 2.17
CodeLlama 2 7B 70.92 84.52 0.75 67.47 5.17 63.61 1.55 72.24 2.77 66.67 2.66 71.03 2.02
LLaMAX Llama 2 7B 67.49 77.89 0.86 61.45 5.37 71.01 1.47 65.40 2.94 55.56 2.80 73.61 1.97
LLaMAX Llama 2 7B Alpaca 69.86 82.75 0.78 66.27 5.22 72.37 1.44 66.16 2.92 59.37 2.77 72.22 2.00
MaLA-500 Llama 2 10B v1 65.89 83.66 0.77 63.86 5.31 60.17 1.58 69.20 2.85 61.90 2.74 56.55 2.21
MaLA-500 Llama 2 10B v2 65.89 83.66 0.77 63.86 5.31 60.17 1.58 69.20 2.85 61.90 2.74 56.55 2.21
YaYi Llama 2 7B 74.50 88.52 0.66 71.08 5.01 72.26 1.45 74.14 2.70 71.75 2.54 69.25 2.06
TowerBase Llama 2 7B 74.29 87.14 0.69 74.70 4.80 69.45 1.49 75.67 2.65 64.76 2.70 74.01 1.96
TowerInstruct Llama 2 7B 74.00 86.28 0.71 77.11 4.64 68.40 1.50 77.19 2.59 66.35 2.67 68.65 2.07
Occiglot Mistral 7B v0.1 74.61 86.54 0.71 79.52 4.46 66.01 1.53 71.86 2.78 66.35 2.67 77.38 1.87
Occiglot Mistral 7B v0.1 Instruct 72.93 85.89 0.72 74.70 4.80 64.55 1.55 70.72 2.81 64.76 2.70 76.98 1.88
BLOOM 7B 70.13 82.24 0.79 71.08 5.01 58.81 1.59 76.81 2.61 57.46 2.79 74.40 1.95
BLOOMZ 7B 67.95 83.40 0.77 73.49 4.87 57.56 1.60 67.30 2.90 54.92 2.81 71.03 2.02
mGPT 59.69 62.67 1.00 59.04 5.43 53.49 1.61 57.41 3.05 58.10 2.78 67.46 2.09
mGPT 13B 63.59 70.62 0.94 63.86 5.31 57.77 1.60 61.22 3.01 60.00 2.76 68.06 2.08
YaYi 7B 69.79 84.04 0.76 75.90 4.72 58.08 1.59 72.24 2.77 55.87 2.80 72.62 1.99
Aya 23 8B 65.84 75.31 0.89 62.65 5.34 65.07 1.54 63.50 2.97 60.63 2.76 67.86 2.08
Aya expanse 8B 71.94 78.58 0.85 75.90 4.72 71.43 1.46 70.34 2.82 64.76 2.70 70.63 2.03
Llama 3 8B 76.84 86.80 0.70 71.08 5.01 75.29 1.39 79.85 2.48 71.43 2.55 76.59 1.89
Llama 3.1 8B 75.52 87.57 0.68 66.27 5.22 75.29 1.39 80.61 2.44 67.62 2.64 75.79 1.91
Gemma 7B 77.41 88.00 0.67 73.49 4.87 75.39 1.39 78.33 2.55 72.06 2.53 77.18 1.87
Gemma 2 9B 80.07 89.42 0.64 75.90 4.72 79.77 1.30 82.89 2.33 74.29 2.47 78.17 1.84
Qwen 2 7B 76.44 86.32 0.71 69.88 5.07 68.82 1.50 78.71 2.53 71.75 2.54 83.13 1.67
Qwen 1.5 7B 72.59 83.31 0.77 69.88 5.07 65.90 1.53 70.34 2.82 66.35 2.67 79.76 1.79
EMMA-500 Llama 2 7B 72.80 82.45 0.79 72.29 4.94 71.64 1.46 69.20 2.85 69.21 2.61 72.02 2.00
@@
Table 24: 0-shot results (ACC %) on XNLI in all languages.
@@
Llama 2 7B 40.19 35.42 0.96 42.65 0.99 47.11 1.00 36.67 0.97 55.30 1.00 40.52 0.98 50.08 1.00 37.71 0.97 42.37 0.99 34.94 0.96 36.35 0.96 37.27 0.97 33.61 0.95 36.63 0.97 36.18 0.96
Llama 2 7B Chat 38.58 34.42 0.95 37.07 0.97 43.09 0.99 38.15 0.97 50.24 1.00 39.44 0.98 44.82 1.00 35.78 0.96 42.09 0.99 34.22 0.95 33.49 0.95 36.95 0.97 33.90 0.95 38.11 0.97 36.95 0.97
CodeLlama 2 7B 40.19 33.41 0.95 37.75 0.97 47.23 1.00 37.63 0.97 54.78 1.00 44.38 1.00 49.20 1.00 35.94 0.96 46.06 1.00 33.29 0.94 35.02 0.96 38.59 0.98 33.25 0.94 40.40 0.98 35.94 0.96
LLaMAX Llama 2 7B 44.27 33.78 0.95 46.83 1.00 48.96 1.00 42.57 0.99 54.90 1.00 47.59 1.00 47.79 1.00 45.50 1.00 45.54 1.00 43.05 0.99 41.85 0.99 43.29 0.99 44.18 1.00 43.86 0.99 34.38 0.95
LLaMAX Llama 2 7B Alpaca 45.09 34.42 0.95 46.39 1.00 49.76 1.00 43.41 0.99 58.11 0.99 48.96 1.00 51.97 1.00 45.62 1.00 46.27 1.00 43.57 0.99 40.80 0.99 43.86 0.99 44.30 1.00 43.09 0.99 35.78 0.96
MaLA-500 Llama 2 10B v1 38.11 35.94 0.96 41.20 0.99 47.51 1.00 34.46 0.95 56.18 0.99 34.10 0.95 47.59 1.00 33.65 0.95 33.94 0.95 35.22 0.96 33.69 0.95 33.82 0.95 35.02 0.96 36.02 0.96 33.25 0.94
MaLA-500 Llama 2 10B v2 38.11 35.94 0.96 41.20 0.99 47.51 1.00 34.46 0.95 56.18 0.99 34.10 0.95 47.59 1.00 33.65 0.95 33.94 0.95 35.22 0.96 33.69 0.95 33.82 0.95 35.02 0.96 36.02 0.96 33.25 0.94
YaYi Llama 2 7B 41.28 34.14 0.95 42.61 0.99 48.84 1.00 37.35 0.97 56.47 0.99 45.78 1.00 51.04 1.00 39.48 0.98 46.27 1.00 35.70 0.96 35.62 0.96 39.36 0.98 33.49 0.95 37.51 0.97 35.54 0.96
TowerBase Llama 2 7B 39.84 33.90 0.95 41.37 0.99 47.87 1.00 35.26 0.96 56.35 0.99 41.69 0.99 49.44 1.00 34.54 0.95 45.94 1.00 35.02 0.96 34.78 0.95 35.74 0.96 33.37 0.95 37.19 0.97 35.22 0.96
TowerInstruct Llama 2 7B 40.36 33.65 0.95 42.93 0.99 48.84 1.00 34.98 0.96 56.95 0.99 46.51 1.00 46.43 1.00 34.74 0.95 46.27 1.00 33.94 0.95 33.90 0.95 37.87 0.97 33.53 0.95 37.35 0.97 37.47 0.97
Occiglot Mistral 7B v0.1 42.35 33.86 0.95 41.37 0.99 51.77 1.00 37.71 0.97 55.86 1.00 51.65 1.00 51.93 1.00 35.74 0.96 47.63 1.00 34.70 0.95 37.39 0.97 43.01 0.99 33.49 0.95 38.63 0.98 40.56 0.98
Occiglot Mistral 7B v0.1 Instruct 40.81 34.38 0.95 38.84 0.98 50.84 1.00 40.00 0.98 55.66 1.00 48.63 1.00 51.69 1.00 34.30 0.95 40.04 0.98 33.17 0.94 36.35 0.96 37.99 0.97 34.06 0.95 37.59 0.97 38.63 0.98
BLOOM 7B 41.60 33.85 0.67 39.92 0.69 39.78 0.69 35.37 0.68 53.97 0.70 48.82 0.71 49.80 0.71 46.51 0.70 43.03 0.70 37.88 0.69 35.05 0.67 35.09 0.67 42.20 0.70 47.39 0.71 35.35 0.68
BLOOMZ 7B 37.13 32.69 0.94 34.02 0.95 41.69 0.99 35.82 0.96 46.87 1.00 36.02 0.96 43.05 0.99 40.40 0.98 37.47 0.97 33.61 0.95 33.09 0.94 33.73 0.95 36.83 0.97 36.67 0.97 35.02 0.96
mGPT 40.51 33.82 0.95 41.73 0.99 44.78 1.00 35.42 0.96 49.36 1.00 42.81 0.99 44.06 1.00 41.33 0.99 43.21 0.99 41.89 0.99 35.78 0.96 40.12 0.98 34.22 0.95 45.50 1.00 33.61 0.95
mGPT 13B 41.59 33.37 0.95 44.82 1.00 46.47 1.00 37.71 0.97 49.68 1.00 44.50 1.00 45.62 1.00 43.05 0.99 44.86 1.00 40.32 0.98 38.67 0.98 38.84 0.98 37.47 0.97 44.74 1.00 33.69 0.95
YaYi 7B 39.87 39.80 0.98 35.78 0.96 42.61 0.99 36.47 0.96 50.52 1.00 47.71 1.00 48.19 1.00 40.04 0.98 39.12 0.98 34.06 0.95 34.34 0.95 33.17 0.94 37.07 0.97 44.18 1.00 34.94 0.96
Aya 23 8B 43.12 33.90 0.95 39.36 0.98 49.36 1.00 41.12 0.99 51.57 1.00 50.36 1.00 51.16 1.00 46.67 1.00 48.67 1.00 34.50 0.95 35.90 0.96 48.39 1.00 33.61 0.95 42.25 0.99 39.96 0.98
Aya expanse 8B 45.56 34.10 0.95 41.89 0.99 51.04 1.00 42.17 0.99 53.86 1.00 47.35 1.00 53.65 1.00 48.31 1.00 51.00 1.00 37.03 0.97 41.04 0.99 48.84 1.00 37.07 0.97 50.12 1.00 45.98 1.00
Llama 3 8B 44.97 33.65 0.95 45.34 1.00 50.48 1.00 39.28 0.98 55.02 1.00 49.52 1.00 50.56 1.00 47.55 1.00 49.16 1.00 38.92 0.98 46.27 1.00 48.23 1.00 33.49 0.95 49.00 1.00 38.15 0.97
Llama 3.1 8B 45.62 33.86 0.95 45.58 1.00 51.45 1.00 38.96 0.98 55.22 1.00 50.28 1.00 51.77 1.00 49.40 1.00 49.16 1.00 39.36 0.98 48.07 1.00 49.32 1.00 35.06 0.96 47.11 1.00 39.76 0.98
Gemma 7B 42.58 33.49 0.95 43.49 0.99 48.63 1.00 38.11 0.97 52.05 1.00 44.14 1.00 49.76 1.00 44.34 1.00 47.39 1.00 40.64 0.98 37.95 0.97 43.25 0.99 35.42 0.96 43.29 0.99 36.67 0.97
Gemma 2 9B 46.74 34.18 0.95 49.52 1.00 51.37 1.00 43.25 0.99 53.45 1.00 51.41 1.00 52.29 1.00 47.31 1.00 49.56 1.00 45.58 1.00 49.88 1.00 50.72 1.00 44.02 1.00 45.70 1.00 32.93 0.94
Qwen 2 7B 42.77 33.69 0.95 45.46 1.00 48.19 1.00 36.83 0.97 54.26 1.00 47.23 1.00 51.41 1.00 44.98 1.00 47.39 1.00 37.31 0.97 38.27 0.97 43.53 0.99 34.02 0.95 43.57 0.99 35.38 0.96
Qwen 1.5 7B 39.47 34.34 0.95 40.92 0.99 42.77 0.99 36.18 0.96 49.08 1.00 37.87 0.97 43.13 0.99 38.15 0.97 38.88 0.98 35.42 0.96 44.22 1.00 38.84 0.98 33.86 0.95 44.38 1.00 33.98 0.95
EMMA-500 Llama 2 7B 45.14 34.78 0.95 46.27 1.00 47.07 1.00 45.86 1.00 53.78 1.00 47.07 1.00 46.87 1.00 47.59 1.00 46.55 1.00 46.18 1.00 41.97 0.99 44.86 1.00 45.98 1.00 47.03 1.00 35.22 0.96
@@
Tables 25 and 26 show 3-shot results on MGSM by direct and Chain-of-Thought prompting respectively. All the scores are obtained by flexible matching.
@@
Table 25: 3-shot results (ACC %) on MGSM in all languages by direct prompting and flexible matching.